OpenStudy (anonymous):

Find the common solution. 5x + 10y = 18 x = y - 6 ( ?,? )

OpenStudy (anonymous):

5(y-6)+10y=18 5y-30+10y=18 15y-30=18 15y=48 y=48/15.

OpenStudy (anonymous):

then plug that in to get x.